Agenda/Minutes on 2021/1/8
CRY+VIS Interface Meeting
Participants: Takahashi, Hirata, Sato, Yano, Aso, Miyo, Ushiba, Tomaru, Ikeda, Kimura, Hagiwara
- Schedule for IX: Connection of ITMX was scheduled on 2/9.
- Schedule for IY: Cooling of payload is scheduled from 1/12.
- Schedule for EX: Let's discuss in the next time.
- Corrugate tube fixing jig: Re-measurement at the beginning of March in IYV is planed.
- Commissioning in vacuum:
- Acceptance check items:
- Jig for FC painting: Under manufacturing.
- PF actuator: by the end of January.
Participants: Takahashi, Tanaka, Fabian, Sato, Hirata, Aso, Miyo, Washimi
- Remedy work in IXV.
- Replaced three blades for TF.
- Mounted the ballast weight on the top plate of TF keystone.
- Assembled and tested the limit switches for the yaw and vertical FR.
- Assembled the geophones into the pods.
Week of the 4th of January
- In order to answer the question whether we should move the assembly frame next to SR3 as a back up plan,
- We opened SRM chamber -X door, next to where the frame is now and
- We measured the width of the clear space: 66 cm, which is enough.
- We should ask Mirapro to move the frame, but
- We need to move the rack with electronics and the cables and we need to allocate time to do it.
- The IP stepper motor shaft pitch was measured to be 1.0 mm/thread.
- The IP stepper motor shaft pitch was measured to be 0.8 mm/thread.
- We received special screws, we installed the IP LVDTs and calibrated them: I have not checked the data but we might need to optimize the LVDT cards and repeat the calibration. I need to think about this issue more carefully.
- Before removing IP immobilizing jig we measured the output of the LVDTs in order to recover the aligned position.
- We replaced the heavy F0 top plate of the keystone assembly by a lighter one with ballast mass.
- We installed the F0 yaw stepper motor with limit switches. Test pending.
- We installed the F0 fishing rod limit switches. Test pending.
We checked the flanges we could use for the cables of a thermistor. See details in klog 15700.
- If travel for Kamioka was limited, remedy work in PR3 would be postponed to February.
- Made the ver.2 LVDT combiner circuit. Preliminary result showed enough attenuation. The circuit board will be sent to Kamioka in the next week.
- Type-B remedy work is on going.
- Remedy work in IXV was started. Type-A team will go to installation in EYV parallelly.
- Procedure documents.
- Criteria in the acceptance test.
The week from 1/11
- Takahashi: -1/15
- Sato: -1/15
- Hirata: -1/15
- Yano: -1/15
The week from 1/18
- Takahashi: ?
- Sato: 1/18-22
- Hirata: ?
- Yano: 1/18-22
- On 2020/1/15(Fri)